Print

Print


 Ticket Number: HEA-NOC/19980807-01         Ticket Status: CLOSE
 Ticket Type:   Scheduled                   Ticket Source: HEA-NOC
 Ticket Scope:  Site                        Site/Line:     HEAnet <==> UCD
 Ticket Owner:  [log in to unmask]             Problem Fixer: HEA-NOC
 Ticket Opened:                             Problem Start: 19980810  18:00 UTC+1
 Ticket Closed: 19980810  18:22:31 UTC      Problem End:   19980810  18:20 UTC+1


Problem Description:

UCD was moved to the new HEAnet backbone router Moskva.

Affected:

Outage of approximately one minute during which time access to and
from UCD was blocked.

Actions:

UCD added 137.43.0.0 to the list of OSPF routes and reconfigured
Hazel's serial interface. NOC proceeded to switch the connection
and ensured all routes were being advertised correctly.

Time to Fix:

Approx. 1 minute; switch completed at 18:20.